OpenStudy (anonymous):

Simplify. csc^2x-1 / cot^2x

OpenStudy (anonymous):

\[\frac{ \csc^2x-1 }{ \cot^2x }\]

OpenStudy (anonymous):

@mathstudent55

OpenStudy (timbo5lice):

csc^2(x-1)tan^2(x)

OpenStudy (michele_laino):

we have to use these identities: \[\large \begin{gathered} \csc x = \frac{1}{{\sin x}} \hfill \\ \cot x = \frac{{\cos x}}{{\sin x}} \hfill \\ \end{gathered} \]

OpenStudy (timbo5lice):

@jaysabelle hope this helps :3

OpenStudy (anonymous):

Wait, I don't get it . . . Ugh, I hate trigonometry :/

OpenStudy (timbo5lice):

\[\csc^2(x-1)\tan^2(x)\]

OpenStudy (timbo5lice):

this is that correct term yeah its hard but mathway.com can help a lot ^_^

OpenStudy (michele_laino):

hint: \[\large \frac{{{{\left( {\csc x} \right)}^2}}}{{{{\left( {\cot x} \right)}^2}}} = \frac{{\frac{1}{{{{\left( {\sin x} \right)}^2}}}}}{{\frac{{{{\left( {\cos x} \right)}^2}}}{{{{\left( {\sin x} \right)}^2}}}}} = \frac{1}{{{{\left( {\sin x} \right)}^2}}} \times \frac{{{{\left( {\sin x} \right)}^2}}}{{{{\left( {\cos x} \right)}^2}}} = ...?\]

OpenStudy (anonymous):

@timbo5lice So, you're saying that \[\csc^2(x-1)\tan^2(x)\] is the most simplified it cab be?

OpenStudy (timbo5lice):

yes ^_^

OpenStudy (michele_laino):

oops.. \[\large \frac{{{{\left( {\csc x} \right)}^2} - 1}}{{{{\left( {\cot x} \right)}^2}}} = \frac{{\frac{1}{{{{\left( {\sin x} \right)}^2}}} - 1}}{{\frac{{{{\left( {\cos x} \right)}^2}}}{{{{\left( {\sin x} \right)}^2}}}}} = \frac{{1 - {{\left( {\sin x} \right)}^2}}}{{{{\left( {\sin x} \right)}^2}}} \times \frac{{{{\left( {\sin x} \right)}^2}}}{{{{\left( {\cos x} \right)}^2}}} = ...?\]

OpenStudy (anonymous):

@Michele_Laino could you help me understand more of it? Trig really is my weakest subject.

OpenStudy (michele_laino):

can you continue from my last step?

OpenStudy (anonymous):

No, I don't know how to go about it

OpenStudy (timbo5lice):

try mathway.com its a calculator for this stuff (:

OpenStudy (michele_laino):

hint: